Week 48


Machine coordinators Roger Bailey and Gianluigi Arduini

Slides shown at meetings





No beam











































Status Monday morning


Both beams captured

Instrumentation commissioning well advanced

Beam 1

        RF phase adjusted and phase lock loop on

        Lifetime 5-10 hours

        Orbit Q Q dispersion all look good

        Coupling and beta-beat need correction

Beam 2

        RF phase adjusted

        Lifetime 3 hours

        Orbit Q Q coupling measured and corrected


Outstanding issues

        Pre-cycling circuits still causing problems

        Still teething problems with new QPS system which is active on sector 12

        Beam dump synchronisation (with RF re-phasing)

        No beam Tuesday afternoon (Radiation survey in the injectors)

        Access in LHC before and during this

o   QPS software upgrade on 600A systems all around the machine

o   Try to limit access beyond this (some hope!)




Beam 2 brought to the same level as Beam 1 (Monday morning)

        Both beams with RF phase locked

        Orbit, Q, Q, coupling OK to first order

        Lifetimes 10 hours

        Dispersion excellent

        Beta-beat measured

Two beam runs (Monday afternoon)

        Beams steered with BPMs interpolated to IP, successively IP1, IP2, IP8, IP5

        Collisions in ATLAS, ALICE, LHCB and CMS

First test ramp (Monday midnight)

        Beam to around 540GeV

        Lost due to tune drifting

        Collimators did their job


Scheduled access given Tuesday morning for machine wide QPS modifications

Many other interventions for a variety of things

Difficult recovery overnight Tuesday/Wednesday (many subsequent interventions needed for QPS)

Operation with beam then further held up by

        a problem in LHC cryogenics in Sector 12 (Wednesday 09-19)

        a vacuum leak in the PS (Thursday 02-16)

Just 2 hours of beam between Tuesday morning and Thursday afternoon

However significant progress made with LHC superconducting circuit cycling

LHC precycle now working

This allows operation to re-establish a reproducible machine magnetically


Restarted with beam Thursday evening 17.00

Beam 2 injected without problems. Tunes come back to how they were yesterday.

Beam 1 problems to inject

        Uncertainty of TCDQ position

        RF synchronisation problems

        Beam 1 injected around 22.30

Energy matching done using beam 2, then verified with beam 1

Frequency changed 106Hz and SPS energy increased 250MeV, then TI2 and TI8 trimmed to match

Looks good, as can be seen on first turn trajectory Beam 2, orbits Beam 1 and Beam 2

Operation with beam then stopped again, due to a problem with LHC cryogenics in Sector 34


Cryo back at 18.00 Friday but problems to get all 600A on due to QPS issues ...

Finally all (needed) circuits OK at 23.00

Struggled with the injection interlock system. Beam in around 01.30


        Re-established conditions for beams 1 and 2

        Kick response beam 2 (sectors 12 23 34 45 done)

        Beam dumps due to beam loss when starting S56

        PS off from 04-07


Beam dump studies started at 08.00 Saturday

        Synchronisation checked with beam by scanning RF bucket through sweeps for B1 and B2

        Protection devices and aperture scans

o   Aperture scanned for B1 extraction channel

o   Apertures look similar to those obtained for B2 on Monday

o   Physical apertures seem OK on both planes

        Verification of the protection in case of asynchronous dump


Operations measurements and tuning in the afternoon/evening


o   Reference orbit B1 and B2 using common correctors

        Tune, coupling and chromaticity measurement and correction

o   C_ down to ~0.005 (Beam 1)/0.0025 (Beam 2) after correction


o   Measured values from +- 100, 50,0 trims off ref
Beam 1
QH=+5 - QV=+7
Beam 2
QH=+9 - QH=+8

        Undulator RU.L4 ON at 50 A

o   no effect on orbit/tune (Beam 2) but not sufficient to see synch light HW commissioning to be continued


BI systems commissioning in parallel through the afternoon/evening

        phased in horizontal and vertical MKQA tune kicker for B1 and B2

        measured and corrected PLL beam transfer function for B1 and B2 system (B2 requires some polarity/mapping checks)

o   tested continuous radial modulation via the orbit feedback controller and verified switch on/off procedures, frequency and amplitudes behaviour

o   performed initial PLL lock studies with B1, initial results look very promising, further assessment w.r.t. nominal PLL performance (bandwidth/sensitivity to resolve dp/p changes in the order of 1e-5) is needed

o   performed semi-automatic chromaticity measurements and preliminary tune modulation/de-modulation studies in view of continuous Q' tracking -> some FESA/SW issues have been identified that need to be addressed

        Not yet ready for tune/chromaticity feedback


Beta-beating studies overnight

        The beta-beating in beam2 was very different to the beta-beating measured before the ramp (most likely due to the lack of precycling then).

        For Beam1 the initial beta beating agrees with the measured after the ramp (Friday 00:00 am), proving no changes in time with proper pre-cycling.

        The first local error was found and successfully corrected in IP2 (MQXB2.R2 + 7e-5 and MQXB2.L2 -5e-5).

o   A reduction of the peak vertical beta-beating in beam 1 of 10% was observed.

o   No change was observed in the peak b-beat in beam 2 and the horizontal of beam1.

        We removed all the trims and re-measured the initial conditions without observing any change (no hysteresis effects).

        Another study in IP7 was started but due to lack of time and frequent beam absence was not conclusive.


Collimator set up Sunday


o   All IR3, IR7 to nominal settings (primary aperture about 5.7 sigma) ==> multi-stage collimation is set up
Dump protection IR6 set up
TCT's in IR1 cross-checked: rough setting confirmed
Other TCT's not touched: at +-15 mm around +-15-25 sigma


o   All IR3, IR7 to nominal settings (primary aperture about 5.7 sigma) ==> mutli-stage collimation is set up
Dump protection IR6 set up
TCT's not touched: at +-15 mm around +-15-25 sigma

        Additional tests:

o   Check on interlock threshold for tungsten collimators (see logbook entry)
Movement and beam loss response check for vertical Roman Pots, beam1


Summary of the (shortened) MPS tests

        Tested the reaction time of the FMCMs on RD1.LR1, RD1.LR5, RD34.LR3, RQ4.LR7 by switching off the PCs (to avoid a WIC trigger that is even faster than the FMCM in such a case). In all cases the FMCM dumped the beam before any BLM triggered.

        Checked the reaction of the BLMs by trimming RD1.LR1 and RD1.LR5 by a small step (also to be able to analyze with PM data). In those cases the BLMs in IR7 triggered the dump.

        Tested the SIS interlock on orbit correctors using a long bump over the arc R2-L3. SIS triggered the dump while the bump was ramping up.

        All MPS tests were successful - excellent performance of the FMCMs.


Second LHC ramp

        Beam to 1.04TeV

        Lost due to tune drifting

        Orbit difference wrt reference and orbit evolution during the ramp


Third LHC ramp

        Tunes for beam 1 were different when we returned to injection energy (no precycle)

        Trimmed and incorporated

        Ramped 2 beams, all the way to 1.18TeV !!